# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 3. Pluri-harmonic metrics of Higgs sheaves over the regular loci
We introduce some basic definitions to fix the conventions as follows. The Hitchin-Simpson connection ([97]) is defined by
$\bar{\partial}_{\theta}:=\bar{\partial}_{E}+\theta,\quad D_{H,\theta}^{1,0}:=\partial_{H}+\theta^{*H},\quad D_{H,\theta}=\bar{\partial}_{\theta}+D_{H,\theta}^{1,0},$
@@
Let $X$ be a compact Kähler klt space of dimension $n$. Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is semistable with respect to some Kähler forms $\omega_{0},\cdots,\omega_{n-2}$, then the following orbifold Bogomolov-Gieseker inequality holds.
@@
If $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{pHiggs}_{X_{\mathrm{reg}}}$, then $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ admits a pluri-harmonic metric $H_{X\setminus\Sigma}$ outside an analytic subspace $\Sigma\subsetneq X$ with $\mathrm{codim}_{X}\Sigma\geq 2$ and $X_{\text{Sing}}\subseteq\Sigma$ such that $\mathcal{E}_{X_{\mathrm{reg}}}|_{X\setminus\Sigma}$ is locally free.
@@
When $X$ is smooth and $\mathcal{E}_{X_{\mathrm{reg}}}$, Theorem 3.2 builds on the Donaldson-Uhlenbeck-Yau theorem [24, 104] and its generalization to Higgs bundles [97]; further developments appears in [2, 3, 71, 84, 85, 59, 73, 7, 10, 56, 57], etc. We refer to the introduction of [90] for an overview of developments on the orbifold Bogomolov-Gieseker inequality.
@@
    The construction of an orbifold Higgs field of the pull-back of a reflexive Higgs sheaf via Kebekus-Schnell’s work [63] on pulling back reflexive differentials (cf. Proposition 3.4).
@@
Once these ingredients are in place, the subsequent proof at this stage can be reduced to study the limiting behavior of the Hermitian-Yang-Mills flow over a non-compact balanced manifold $(X\setminus\Sigma,\omega)$ and the strategy in [73, Section 4] originated from [97] remains valid.
3.1.1. Pull-back of the Higgs field over orbifold modifications
We first review the construction in [110, Section 5.1].
@@ 3.1.1. Pull-back of the Higgs field over orbifold modifications
Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is a reflexive Higgs sheaf on the regular locus of a compact complex klt space $X$ and $\omega_{0},\cdots,\omega_{n-2}$ be Kähler forms, there exists a projective bimeromorphism $f:Y\rightarrow X$ from a compact Kähler space $Y$ with only quotient singularities to $X$ with the following data.
@@
    An complex effective orbifold structure $Y_{\mathrm{orb}}:=\{U_{\mu},G_{\mu},\pi_{\mu}\}$ of $Y$.
@@
    An analytic subspace $\Sigma\subset X$ of codimension at least $2$ such that $f_{\mu}:=f\circ\pi_{\mu}$ is étale outside $f_{\mu}^{-1}(\Sigma)$ for each $i$, $X$ is smooth and $\mathcal{E}_{X_{\mathrm{reg}}}$ is locally free outside $\Sigma$.
@@ A Higgs orbi-bundle $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ over $Y_{\mathrm{orb}}$ with
    (3.2) $c_{1}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot f^{*}\gamma=\widehat{c}_{1}(\mathcal{E}_{X})\cdot\gamma=c_{1}(\mathcal{F})\cdot\gamma,\ \forall\gamma\in H_{2n-2}(X,\mathbb{R})$
@@
    (3.4) $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})=\{f_{\mu}^{*}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\}$
    outside $f_{\mathrm{orb}}^{-1}(\Sigma):=\{f_{\mu}^{-1}(\Sigma)\}$. The equality 3.3 also holds for $\widehat{c}_{1}^{2}(\mathcal{E}_{X})$.
@@
Let $h:W\rightarrow X$ be a partial orbifold resolution given by [67] and $W_{\mathrm{orb}}:=\{(V_{\mu},G_{\mu},\pi_{\mu})\}$ be the standard orbifold structure of $W$. Set $\mathcal{E}_{\mathrm{orb}}=h_{\mathrm{orb}}^{[*]}\mathcal{E}_{X}:=\{\big((h\circ\pi_{\mu})^{*}\mathcal{E}_{X}\big)^{\vee\vee}\}.$ According to Lemma 2.2, there exists a projective orbifold morphism $g_{\mathrm{orb}}:Y_{\mathrm{orb}}\rightarrow W_{\mathrm{orb}}$ from a compact complex effextive orbifold $Y_{\mathrm{orb}}:=\{U_{\mu},G_{\mu},\pi_{\mu}\}$ to $W_{\mathrm{orb}}$ such that $E_{\mathrm{orb}}:=g^{T}_{\mathrm{orb}}\mathcal{E}_{\mathrm{orb}}=\{g_{\mu}^{*}\mathcal{E}/(\mathrm{torsion})\}$ is a vector orbi-bundle. Let $g$ be the induced morphism of $g_{\mathrm{orb}}$ from the quotient space $Y$ of $Y_{\mathrm{orb}}$ to $W$, then $f:=h\circ g$ is a projective bimeromorphism.
Building on Lemma 2.10, we conclude that there exists a Higgs field $\theta_{\mathrm{orb}}=\{\theta_{i}\}$ of $E_{\mathrm{orb}}$ such that $\theta_{i}|_{f_{i}^{-1}(X_{\mathrm{reg}})}$ is the pull-back of $\theta$ via $f_{i}$. We summarize the dimensional properties by the constructions:
@@
    There exists an analytic subspace $Z\subset X$ of codimension at least $3$ such that $g$ is biholomorphic outside $h^{-1}(Z)$.
@@
    $W_{\mathrm{orb}}$ is standard, i.e. $\pi_{i}$ is finite and quasi-étale for each $i$.
@@
    $g_{i}$ is biholomorphic outside the preimage of the non-locally free locus $S_{i}$ of $\mathcal{E}_{i}$, where $S_{i}$ has codimension at least $3$ due to the reflexivity of $\mathcal{E}_{i}$.
Combining Lemma 2.18 and 2.23, the proof is complete by choosing $\Sigma=Z\cup X_{sing}\cup_{i}h_{i}(S_{i})$. ∎
@@ 3.1.2. Bando-Siu’s methods
Suppose that $\omega_{0},\cdots,\omega_{n-2}$ are Kähler forms on $X$. Recall that $Y_{\mathrm{orb}}$ admits an orbifold Kähler form $\omega_{Y_{\mathrm{orb}}}=\{\omega_{\mu}\}$ by Proposition 2.19. Set
(3.5) $\omega_{\mathrm{orb},k,\epsilon}:=f_{\mathrm{orb}}^{\ast}\omega_{k}+\epsilon\omega_{Y_{\mathrm{orb}}}=\{f_{\mu}^{*}\omega_{k}+\epsilon\omega_{\mu}\}$
for all $0\leq k\leq n-2$, where $f_{\mathrm{orb}}^{*}\omega_{k}:=\{f_{\mu}^{*}\omega_{k}\}$. Recall that for each $\epsilon$, there exists a smooth orbifold Hermitian metric $\omega_{\mathrm{orb},\epsilon}$ (see e.g. [80, Pages 279]) such that
@@
(3.7) $\int_{Y_{\mathrm{orb}}}(|\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{\hat{H}}|_{\hat{H}}+|\theta|_{\hat{H},\omega_{\mathrm{orb},\epsilon}}^{2})\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!}\leq\hat{C}_{0},$
for all $\epsilon$. Consider the Hermtian-Yang-Mills flow on the Higgs orbi-bundle $(E_{\mathrm{orb}},\bar{\partial}_{E_{\mathrm{orb}}},\theta_{\mathrm{orb}})$ with the fixed initial metric $\hat{H}_{\mathrm{orb}}:=\{\hat{H}_{\mu}\}$ and with respect to the Hermitian metric $\omega_{\epsilon}$,
@@
where $\lambda_{\epsilon}=\frac{2\pi}{\operatorname{Vol}(Y_{\mathrm{orb}},\omega_{\mathrm{orb},\epsilon})}\mu_{\omega_{\mathrm{orb},\epsilon}}(E_{\mathrm{orb}})$. Following Simpson’s argument ([97]), we obtain the long time existence and uniqueness of the solution to the above flow (3.8). Actually, the needed estimates are valid on each orbifold chart $U_{i}$, whereby computations agree with those in the manifold case (see e.g. [27] for considerations on the existence of Hermitian-Einstein metrics of stable vector orbi-bundles, and [110, Section 3] for general stable Higgs orbi-bundles).
In Section 3.2, building on the uniform Sobolev inequality (Proposition 3.9) and following Bando-Siu’s argument [3], by choosing a subsequence, we shall prove that $H_{\mathrm{orb},\epsilon}(t)$ converges to smooth metrics $H_{\mathrm{orb}}(t)=\{H_{\mu}(t)\}$ in $C_{loc}^{\infty}(Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma))$-topology as $\epsilon\rightarrow 0$ and thus $H_{\mathrm{orb}}(t)$ can be descend to a long time solution $H(t)$ of the Hermitian-Yang-Mills flow on $(X\setminus\Sigma)\times[0,+\infty)$, i.e. $H(t)$ satisfies:
@@
because $\widehat{H}_{\mu}$ is $G_{\mu}$-invariant and $f_{\mu}$ is étale outside $f_{\mu}^{-1}(\Sigma)$, where $\omega$ is a smooth Hermitian metric on $X$ such that
@@
and $\lambda=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f_{\mathrm{orb}}^{*}\omega)^{n}}\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X}).$
In Section 3.3, we shall study the limiting behavior of the Hermitian-Yang-Mills flow (3.8), which enable us to obtain the following theorem.
@@
In the setting in Proposition 3.4, when $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable, then by choosing a subsequence, we have $H(t)\rightarrow H_{\infty}$ in $C^{\infty}_{loc}(X\setminus\Sigma)$ and
$\sqrt{-1}\Lambda_{\omega}(F_{H_{\infty},\theta_{X_{\mathrm{reg}}}})=\lambda\cdot\operatorname{Id}_{\mathcal{E}_{\mathrm{reg}}}$
@@
To establish the Donaldson-Uhlenbeck-Yau theorem in our context, what remains open is the admissibility of $H_{\infty}$. The primary obstacle is proving a uniform estimate of $\theta_{X_{\mathrm{reg}}}$, a necessary step that is currently not attainable (cf. the smooth case in [73, Section 3]).
@@
For $t_{0}>0$, whether $\|\theta_{X_{\mathrm{reg}}}\|_{H(t),\omega}(t>t_{0})$ are uniformly bounded along the Hermitian-Yang-Mill flow $H(t)$?
@@
Building on Theorem 3.5, the proof of Theorem 3.2 can be complete. It suffices to consider the stable case. We borrow the ideas of [10, 9] using the Hodge index Theorem.
Let $X$ be a compact Kähler klt space of dimension $n$ and $\omega_{0},\cdots,\omega_{n-2}$ be n1n-1 Kähler forms. Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ be a $(\omega_{0},\cdots,\omega_{n-2})$-stable reflexive Higgs sheaf on the regular locus of $X$ with vanishing orbifold Chern numbers (1.3). We adapts the notations in Section 3.1.1 and Section 3.1.1. It follows from (3.2) that
@@
$0$ $=(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\omega_{\epsilon})\wedge\omega_{0,\epsilon}\wedge\cdots\wedge\omega_{n-2,\epsilon}$
$=(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\wedge\omega_{0,\epsilon}\wedge\cdots\wedge\omega_{n-2,\epsilon}$
where $\gamma_{\epsilon}=\frac{\omega_{\epsilon}^{n}}{\omega_{0,\epsilon}^{2}\wedge\cdots\wedge\omega_{n-2,\epsilon}}$ ($\epsilon>0$) is a smooth function on $Y_{\mathrm{orb}}$. Recall the following elementary statement ([102]).
@@
Let $V$ be a complex linear space of dimension $n$. Let $\omega_{0},\cdots,\omega_{n-2}$ be n1n-1 positive $(1,1)$-forms. We define a Hermitian form on $\Lambda^{1,1}V$ by
$Q(\alpha,\sigma):=*(\alpha\wedge\bar{\sigma}\wedge\omega_{1}\wedge\cdots\wedge\omega_{n-2})$
@@
$P^{1,1}:=\{\alpha\in\Lambda^{1,1}V:\alpha\wedge\omega_{0}\wedge\cdots\wedge\omega_{n-2}=0\}.$
@@
By applying Lemma 3.7 to $\sqrt{-1}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}(\sqrt{-1}F_{H_{\mathrm{orb},\epsilon}(t)})}{n}\gamma_{\epsilon}\omega_{0,\epsilon}$ with respect to $\omega_{0,\epsilon},\cdots,\omega_{n-2,\epsilon}$, we conclude that
$\operatorname{tr}\left((F_{H_{\mathrm{orb},\epsilon},\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\wedge(F_{H_{\mathrm{orb},\epsilon},\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\epsilon}\omega_{0,\epsilon})\right)\wedge\Omega_{\epsilon}$
is positive outside $f_{\mathrm{orb}}^{-1}(\Sigma)$ for $0<\epsilon\ll 1$, and note that its descend via $f_{\mathrm{orb}}$ converges to

### v2

@@ 3. Pluri-harmonic metrics on Higgs sheaves over the regular locus
We introduce some basic definitions to fix the conventions as follows. The Hitchin-Simpson connection [100] is defined by
$\bar{\partial}_{\theta}:=\bar{\partial}_{E}+\theta,\quad D_{H,\theta}^{1,0}:=\partial_{H}+\theta^{*H},\quad D_{H,\theta}:=\bar{\partial}_{\theta}+D_{H,\theta}^{1,0},$
@@
Let $X$ be a compact Kähler klt space of dimension $n$. Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is semistable with respect to some Kähler forms $\omega_{0},\cdots,\omega_{n-2}$, then the following orbifold Bogomolov-Gieseker inequality holds:
@@
Moreover, if $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{pHiggs}_{X_{\mathrm{reg}}}$, then $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ admits a pluri-harmonic metric $H_{X\setminus\Sigma}$ on $X\setminus\Sigma$, where $\Sigma\subsetneq X$ is an analytic subspace of codimension at least $2$ containing $X_{sing}$, such that $\mathcal{E}_{X_{\mathrm{reg}}}|_{X\setminus\Sigma}$ is locally free.
@@
When $X$ is smooth and $\mathcal{E}_{X_{\mathrm{reg}}}$ is locally free, Theorem 3.2 builds on the Donaldson-Uhlenbeck-Yau theorem [24, 107] and its generalization to Higgs bundles [100]; further developments appear in [2, 3, 72, 85, 86, 60, 74, 7, 10, 57, 58], etc. We refer to the introduction of [92] for an overview of developments on the orbifold Bogomolov-Gieseker inequality.
@@
    The construction of an orbifold Higgs field of the pullback of a reflexive Higgs sheaf via Kebekus-Schnell’s work [64] on pulling back reflexive differentials (cf. Proposition 3.4).
@@
Once these ingredients are in place, the subsequent proof at this stage can be reduced to studying the limiting behavior of the Hermitian-Yang-Mills flow over a noncompact balanced manifold $(X\setminus\Sigma,\omega)$ and the strategy in [74, Section 4] originated from [100] remains valid.
3.1.1. Pullback of the Higgs field over orbifold modifications
We first review the construction in [113, Section 5.1].
@@ 3.1.1. Pullback of the Higgs field over orbifold modifications
Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is a reflexive Higgs sheaf on the regular locus of a compact complex klt space $X$ and $\omega_{0},\cdots,\omega_{n-2}$ are Kähler forms, then there exists a projective bimeromorphism $f:Y\rightarrow X$ from a compact Kähler space $Y$ with only quotient singularities to $X$ with the following data:
@@
    A complex effective orbifold structure $Y_{\mathrm{orb}}:=\{(U_{\alpha},G_{\alpha},\mu_{\alpha})\}$ of $Y$.
@@
    An analytic subspace $\Sigma\subset X$ of codimension at least $2$ containing $X_{sing}$ such that each $f_{\alpha}:=f\circ\mu_{\alpha}$ is étale outside $f_{\alpha}^{-1}(\Sigma)$ and $\mathcal{E}_{X_{\mathrm{reg}}}$ is locally free outside $\Sigma$.
@@ A Higgs orbi-bundle $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ over $Y_{\mathrm{orb}}$ with
    (3.2) $c_{1}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot f^{*}\sigma=\widehat{c}_{1}(\mathcal{E}_{X})\cdot\sigma,\ \forall\sigma\in H_{2n-2}(X,\mathbb{R})$
@@
    (3.4) $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})=\{f_{\alpha}^{*}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\}$
    outside $f_{\mathrm{orb}}^{-1}(\Sigma):=\{f_{\alpha}^{-1}(\Sigma)\}$. The equality (3.3) also holds for $\widehat{c}_{1}^{2}(\mathcal{E}_{X})$.
@@
Let $h:W\rightarrow X$ be an orbifold modification given by [68] and $W_{\mathrm{orb}}:=\{(V_{\alpha},G_{\alpha},\mu_{\alpha})\}$ be the standard orbifold structure of $W$. Set $\mathcal{E}_{\mathrm{orb}}=h_{\mathrm{orb}}^{[*]}\mathcal{E}_{X}:=\{\big((h\circ\mu_{\alpha})^{*}\mathcal{E}_{X}\big)^{\vee\vee}\}.$ According to Lemma 2.2, there exists a projective orbifold morphism $g_{\mathrm{orb}}:Y_{\mathrm{orb}}\rightarrow W_{\mathrm{orb}}$ from a compact complex effective orbifold $Y_{\mathrm{orb}}:=\{(U_{\alpha},G_{\alpha},\mu_{\alpha})\}$ to $W_{\mathrm{orb}}$ such that $E_{\mathrm{orb}}=\{g_{\alpha}^{*}\mathcal{E}/(\mathrm{torsion})\}$ is a vector orbi-bundle. Let $g$ be the induced morphism from the quotient space $Y$ of $Y_{\mathrm{orb}}$ to $W$, then $f:=h\circ g$ is a projective bimeromorphism.
Building on Lemma 2.10, we conclude that there exists a Higgs field $\theta_{\mathrm{orb}}=\{\theta_{\alpha}\}$ of $E_{\mathrm{orb}}$ such that $\theta_{\alpha}|_{f_{\alpha}^{-1}(X_{\mathrm{reg}})}$ is the pullback of $\theta_{X_{\mathrm{reg}}}$ via $f_{\alpha}$. We summarize the properties resulting from the constructions as follows:
@@
    There exists an analytic subspace $Z\subset X$ of codimension at least $3$ such that $h$ is biholomorphic outside $h^{-1}(Z)$.
@@
    The orbifold structure $W_{\mathrm{orb}}$ is standard, i.e., $\mu_{\alpha}$ is finite and quasi-étale for each $\alpha$.
@@
    The morphism $g_{\alpha}$ is biholomorphic outside the preimage of the non-locally free locus $S_{\alpha}$ of $\mathcal{E}_{\alpha}$, where $S_{\alpha}$ has codimension at least $3$ due to the reflexivity of $\mathcal{E}_{\alpha}$.
Combining Lemma 2.18 and 2.23, the proof is complete by choosing $\Sigma=Z\cup X_{sing}\cup(\cup_{\alpha}h_{\alpha}(S_{\alpha}))$. ∎
@@ 3.1.2. Bando-Siu’s methods
Suppose that $\omega_{0},\cdots,\omega_{n-2}$ are Kähler forms on $X$. By Proposition 2.19, $Y_{\mathrm{orb}}$ admits an orbifold Kähler form $\omega_{Y_{\mathrm{orb}}}=\{\omega_{\alpha}\}$. Set
(3.5) $\omega_{\mathrm{orb},k,\epsilon}:=f_{\mathrm{orb}}^{\ast}\omega_{k}+\epsilon\omega_{Y_{\mathrm{orb}}}=\{f_{\alpha}^{*}\omega_{k}+\epsilon\omega_{\alpha}\}$
for all $0\leq k\leq n-2$, where $f_{\mathrm{orb}}^{*}\omega_{k}:=\{f_{\alpha}^{*}\omega_{k}\}$. Recall that for each $\epsilon$, there exists a smooth orbifold Hermitian metric $\omega_{\mathrm{orb},\epsilon}$ (see e.g. [81, Pages 279]) such that
@@
(3.7) $\int_{Y_{\mathrm{orb}}}(|\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{\hat{H}}|_{\hat{H}}+|\theta|_{\hat{H},\omega_{\mathrm{orb},\epsilon}}^{2})\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{n!}\leq\hat{C}_{0}$
for all $\epsilon$. Consider the Hermitian-Yang-Mills flow on the Higgs orbi-bundle $(E_{\mathrm{orb}},\bar{\partial}_{E_{\mathrm{orb}}},\theta_{\mathrm{orb}})$ with the fixed initial metric $\hat{H}_{\mathrm{orb}}:=\{\hat{H}_{\alpha}\}$ and with respect to the Hermitian metric $\omega_{\mathrm{orb},\epsilon}$,
@@
where $\lambda_{\epsilon}=\frac{2\pi}{\operatorname{Vol}(Y_{\mathrm{orb}},\omega_{\mathrm{orb},\epsilon})}\mu_{\omega_{\mathrm{orb},\epsilon}}(E_{\mathrm{orb}})$. Following Simpson’s argument [100], we obtain the long time existence and uniqueness of the solution to the above flow (3.8). Actually, the needed estimates are valid on each orbifold chart $U_{\alpha}$, where the computations are consistent with those in the manifold case (see e.g. [27] for the existence of Hermitian-Einstein metrics on stable vector orbi-bundles, and [113, Section 3] for the general case of stable Higgs orbi-bundles).
In Section 3.2, building on the uniform Sobolev inequality (Proposition 3.9) and following Bando-Siu’s argument [3], by choosing a subsequence, we shall prove that $H_{\mathrm{orb},\epsilon}(t)$ converges to smooth metric $H_{\mathrm{orb}}(t):=\{H_{\alpha}(t)\}$ in $C_{loc}^{\infty}(Y_{\mathrm{orb}}\setminus f_{\mathrm{orb}}^{-1}(\Sigma))$-topology as $\epsilon\rightarrow 0$ and thus $H_{\mathrm{orb}}(t)$ can descend to a long time solution $H(t)$ of the Hermitian-Yang-Mills flow on $(X\setminus\Sigma)\times[0,+\infty)$, i.e., $H(t)$ satisfies:
@@
because $\widehat{H}_{\alpha}$ is $G_{\alpha}$-invariant and $f_{\alpha}$ is étale outside $f_{\alpha}^{-1}(\Sigma)$, where $\omega$ is a smooth Hermitian metric on $X$ such that
@@
and $\lambda=\frac{2\pi}{\int_{Y_{\mathrm{orb}}}(f_{\mathrm{orb}}^{*}\omega)^{n}}\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})$. Observe that $\{H_{\alpha}(t)\}=\{f_{\alpha}^{*}H(t)\}.$
In Section 3.3, we study the limiting behavior of the Hermitian-Yang-Mills flow (3.8), which enables us to obtain the following theorem.
@@
In the setting of Proposition 3.4, if $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable, then by choosing a subsequence, we have $H(t)\rightarrow H_{\infty}$ in $C^{\infty}_{loc}(X\setminus\Sigma)$ and
$\sqrt{-1}\Lambda_{\omega}(F_{H_{\infty},\theta_{X_{\mathrm{reg}}}})=\lambda\cdot\operatorname{Id}_{\mathcal{E}_{X_{\mathrm{reg}}}}$
@@
To establish the Donaldson-Uhlenbeck-Yau theorem in our context, what remains open is the admissibility of $H_{\infty}$. The primary obstacle is proving a uniform estimate of $\theta_{X_{\mathrm{reg}}}$, a necessary step that is currently not attainable (cf. the smooth case in [74, Section 3]).
@@
For $t_{0}>0$, is $\|\theta_{X_{\mathrm{reg}}}\|_{H(t),\omega}$ uniformly bounded for $t>t_{0}$ along the Hermitian-Yang-Mills flow $H(t)$?
@@
Building on Theorem 3.5, the proof of Theorem 3.2 can be completed. It suffices to consider the stable case. We borrow the ideas of [10, 9] using the Hodge index Theorem.
Let $X$ be a compact Kähler klt space of dimension $n$ and $\omega_{0},\cdots,\omega_{n-2}$ be n1n-1 Kähler forms. Suppose that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is an $(\omega_{0},\cdots,\omega_{n-2})$-stable reflexive Higgs sheaf on the regular locus of $X$ with vanishing orbifold Chern numbers (1.3). We adapt the notations in Section 3.1.1 and Section 3.1.2. It follows from (3.2) that
@@
$0$ $=(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\omega_{\mathrm{orb},\epsilon})\wedge\omega_{\mathrm{orb},0,\epsilon}\wedge\cdots\wedge\omega_{\mathrm{orb},n-2,\epsilon}$
$=(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t)}}{n}\cdot\gamma_{\mathrm{orb},\epsilon}\omega_{\mathrm{orb},0,\epsilon})\wedge\omega_{\mathrm{orb},0,\epsilon}\wedge\cdots\wedge\omega_{\mathrm{orb},n-2,\epsilon},$
where $\gamma_{\mathrm{orb},\epsilon}:=\frac{\omega_{\mathrm{orb},\epsilon}^{n}}{\omega_{\mathrm{orb},0,\epsilon}^{2}\wedge\cdots\wedge\omega_{\mathrm{orb},n-2,\epsilon}}$ ($\epsilon>0$) is a smooth function on $Y_{\mathrm{orb}}$. Recall the following elementary statement [105].
@@
Let $V$ be a complex vector space of dimension $n$. Let $\omega_{0},\cdots,\omega_{n-2}$ be n1n-1 positive $(1,1)$-forms. We define a Hermitian form on $\Lambda^{1,1}V$ by
$Q(\xi,\sigma):=*(\xi\wedge\bar{\sigma}\wedge\omega_{1}\wedge\cdots\wedge\omega_{n-2}),\ \forall\ \xi,\sigma\in\Lambda^{1,1}V,$
@@
$P^{1,1}:=\{\xi\in\Lambda^{1,1}V:\xi\wedge\omega_{0}\wedge\cdots\wedge\omega_{n-2}=0\}.$
@@
By applying Lemma 3.7 to $\sqrt{-1}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\mathrm{orb},\epsilon}}(\sqrt{-1}F_{H_{\mathrm{orb},\epsilon}(t)})}{n}\gamma_{\mathrm{orb},\epsilon}\omega_{\mathrm{orb},0,\epsilon}$ with respect to $\omega_{\mathrm{orb},0,\epsilon},\cdots,\omega_{\mathrm{orb},n-2,\epsilon}$, we conclude that
$\operatorname{tr}\bigg($ $(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}}{n}\cdot\gamma_{\mathrm{orb},\epsilon}\omega_{\mathrm{orb},0,\epsilon})$
$\wedge(F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}-\frac{\Lambda_{\omega_{\mathrm{orb},\epsilon}}F_{H_{\mathrm{orb},\epsilon}(t),\theta_{\mathrm{orb}}}}{n}\cdot\gamma_{\mathrm{orb},\epsilon}\omega_{\mathrm{orb},0,\epsilon})\bigg)\wedge\Omega_{\mathrm{orb},\epsilon}$
is positive outside $f_{\mathrm{orb}}^{-1}(\Sigma)$ for $0<\epsilon\ll 1$, and note that its descent via $f_{\mathrm{orb}}$ converges to